# Barriers for Learning in an Evolving World: Mathematical Understanding of Loss of Plasticity

### v1

@@ A.3 Modular cloning profiles and a composition theorem
Fix a module $M$ with profile $(\mathcal{P}^{\mathrm{in}}_{M},\mathcal{P}^{\mathrm{out}}_{M})$. Extend these interface partitions to a partition of all nodes $V_{M}$ by assigning each internal node of $M$ to the block of its corresponding base-node in the collapsed base graph $\widetilde{G}$ defined in Sec.˜A.2. On $M$, define the (affine) module cloning manifold
@@
i.e., each inter-block weight submatrix is row- and column-equitable (block-wise constant up to redistribution), reusing the notation of Sec.˜A.2. This generalizes the block-constant manifold $\mathcal{M}_{BC}$ by allowing intra-block redistribution while preserving equal in/out block-sums.
@@
(MC3) implies closure under any first-order optimizer whose update is a (possibly stateful) scalar multiple of the local gradient on each parameter and whose internal state is identical across clones at initialization (e.g., SGD, momentum, RMSProp, Adam with tied clone states). Weight decay that acts per-parameter independently may break exact symmetry; see Sec.˜A.2. Dropout violates (MC1)–(MC2) because independent masks destroy blockwise equality in the forward/backward signals.
@@
This is the restriction of Theorem˜2.1 to the subgraph $G_{M}$ with its node partition: row-equitability yields identical forward values within blocks, column-equitability yields identical backward adjoints within blocks, and $d\mathcal{L}/dW_{M}=h\,\delta^{\top}$ is block-wise constant across inter-block submatrices. Tangency of the gradient to $\mathcal{M}_{RE}\cap\mathcal{M}_{CE}$ follows exactly as in Sec.˜A.2. ∎
@@ B.1.3 General Setup and Procedures
For the continual learning and cloning experiments, our experimental grid consisted of approximately 2,000 individual runs (counting each random seed separately). These experiments were executed on a cluster of NVIDIA A100 GPUs, utilizing a heterogeneous mix of 40GB and 80GB memory variants. The total computational cost for these experiments was approximately 10,000 GPU-hours. The bit flipping experiments and additional theory validation experiments were conducted on a more diverse set of hardware, utilizing lower-end computational nodes equipped with NVIDIA RTX 3090, V100, and RTX 2080 GPUs. This heterogeneous setup was sufficient for these less computationally intensive experiments, and the overall compute amounted to under 100 GPU-hours on these nodes. The theoretical validation figures and numerical simulations presented in the theory appendix (Section˜A.1.1) were generated on a MacBook using CPU computation only.

### v2

@@ A.4 Modular cloning profiles and a composition theorem
Fix a module $M$ with profile $(\mathcal{P}^{\mathrm{in}}_{M},\mathcal{P}^{\mathrm{out}}_{M})$. Extend these interface partitions to a partition of all nodes $V_{M}$ by assigning each internal node of $M$ to the block of its corresponding base-node in the collapsed base graph $\widetilde{G}$ defined in Sec.˜A.3. On $M$, define the (affine) module cloning manifold
@@
i.e., each inter-block weight submatrix is row- and column-equitable (block-wise constant up to redistribution), reusing the notation of Sec.˜A.3. This generalizes the block-constant manifold $\mathcal{M}_{BC}$ by allowing intra-block redistribution while preserving equal in/out block-sums.
@@
(MC3) implies closure under any first-order optimizer whose update is a (possibly stateful) scalar multiple of the local gradient on each parameter and whose internal state is identical across clones at initialization (e.g., SGD, momentum, RMSProp, Adam with tied clone states). Weight decay that acts per-parameter independently may break exact symmetry; see Sec.˜A.3. Dropout violates (MC1)–(MC2) because independent masks destroy blockwise equality in the forward/backward signals.
@@
This is the restriction of Theorem˜2.1 to the subgraph $G_{M}$ with its node partition: row-equitability yields identical forward values within blocks, column-equitability yields identical backward adjoints within blocks, and $d\mathcal{L}/dW_{M}=h\,\delta^{\top}$ is block-wise constant across inter-block submatrices. Tangency of the gradient to $\mathcal{M}_{RE}\cap\mathcal{M}_{CE}$ follows exactly as in Sec.˜A.3. ∎
@@ B.1.3 General Setup and Procedures
For the continual learning and cloning experiments, our experimental grid consisted of approximately 2,000 individual runs (counting each random seed separately). These experiments were executed on a cluster of NVIDIA A100 GPUs, utilizing a heterogeneous mix of 40GB and 80GB memory variants. The total computational cost for these experiments was approximately 10,000 GPU-hours. The bit flipping experiments and additional theory validation experiments were conducted on a more diverse set of hardware, utilizing lower-end computational nodes equipped with NVIDIA RTX 3090, V100, and RTX 2080 GPUs. This heterogeneous setup was sufficient for these less computationally intensive experiments, and the overall compute amounted to under 100 GPU-hours on these nodes. The theoretical validation figures and numerical simulations presented in the theory appendix (Section˜A.2) were generated on a MacBook using CPU computation only.